How much work does the electric motor of a Van de Graaff generator do to lift a positive ion (q=e) if the potential of the spherical electrode is 1.8 MV ?
Answer #1

The change in the energy of ion in lifting it by a potential difference of 1.8 MV is

E = e V= 1.602 \times 10^{-19} \times 1.8 \times 10^{6}= 2.884 \times 10^{-13} \,\, J

Now from the work energy theorem, the change in energy of the ion must be equals to the work done by the electric motor on it.

Hence, the electric motor has done a work of 2.884 X 10-13 J in lifting the ion.